# Hongzhi Zang1, Mingjie Wei6,2, Si Xu3, Yongji Wu5, Zhen Guo3, Yuanqing Wang4,3, Hao Lin3, &Liangzhi Shi1, Yuqing Xie1, Zhexuan Xu1, Zhihao Liu7,2, Kang Chen4,2, Wenhao Tang1, &Quanlu Zhang3, Weinan Zhang6, Chao Yu1,2,†,∗, Yu Wang1,†

### v1

@@ 4.1.3 Supported Granularity Combinations for Advantage and Log-Probability
4.2 Design Choices
4.2.1 PPO
Partial Reset Support
In reinforcement learning rollouts, a sub-environment terminates (“done”) when it either reaches the maximum episode length or the task is successfully completed. There are two common strategies for handling such terminations. One is the Fixed Episode Length mode, where environments are only reset after all sub-environments reach the maximum episode length. The other is the Partial Reset mode, where any sub-environment is reset immediately upon termination, without waiting for others. Figure 5(a) and Figure 5(b) illustrate the difference between these two modes. Within RLinf, to enable Fixed Episode Length, please set “ignore_terminations=True”, while to enable Partial Reset, please set “auto_reset=True” and “ignore_terminations=False”.
For PPO, users can flexibly choose between the two modes, which correspond to different optimization objectives. The Fixed Episode Length mode encourages the policy to maintain success until the end of an episode, aligning with the “success_at_end” objective defined in ManiSkill (Tao et al., 2025). In contrast, the Partial Reset mode emphasizes achieving success at least once during a rollout, i.e., the “success_once” objective. When the target is “success_once,” Partial Reset can substantially improve sample efficiency, as successful episodes are immediately reset and new trajectories are collected.
(a) Fixed Episode Length
(b) Partial Reset
(c) Valid Action Mask
Figure 5: Illustration of the three rollout modes. Section 4.2.1 and Section 4.2.2 provide detailed explanations of modes (b) and (c). In the RLinf-VLA framework, modes (a) and (b) are adopted for PPO, while modes (a) and (c) are adopted for GRPO.
Critic Design
To support PPO in VLA, the design of the critic is crucial. Using an entirely separate critic model would be computationally expensive and complicate GPU resource allocation, given the large size of VLA. To address this, we let the actor and critic share most of their parameters. Following the design in RL4VLA (Liu et al., 2025), we attach a lightweight value head (a three-layer MLP) on top of the language model component of VLA to estimate state values efficiently. Specifically, the value head takes the hidden representation $h^{0}$ from the final transformer block (Vaswani et al., 2023) at the position of the first action token and maps it to a scalar value prediction.
Value for Action Chunks
As discussed in Figure 4, two formulations of advantage estimation naturally lead to two corresponding formulations of value estimation for an action chunk $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$:
  1. 1.
    Chunk-level: The critic estimates the value of the entire chunk as a macro-action. Formally, $V:\mathcal{S}\to\mathbb{R}$.
  2. 2.
    Action-level: Each atomic action $a_{t,j}$ is assigned its own value estimate. Since only the initial observation $o_{t}$ is available at the beginning of the chunk, the critic outputs a $C$-dimensional vector, providing one value prediction per step within the chunk. Formally, $V:\mathcal{S}\to\mathbb{R}^{C}$.
Our framework supports both formulations, controlled by the parameter value_type. By default, the configuration sets value_type to match reward_type, ensuring consistency between reward aggregation and value estimation. We report results for both options using the OpenVLA-OFT model, and in general, the action-level formulation yields better performance.
4.2.2 GRPO
Group Design
The core of the GRPO algorithm lies in how groups are defined. In our framework, two trajectories are assigned to the same group if and only if they correspond to the same task and share the same initial state. This design mirrors the LLM setting, where multiple responses belong to the same group if and only if they are generated in response to the same prompt.
Valid Action Mask
During rollout, each environment runs for at most “max_episode_steps” steps. However, many tasks can be successfully completed earlier. This leads to two alternative strategies for policy optimization: (1) using the entire trajectory regardless of when success occurs, or (2) considering only the steps before task completion. Consistent with the Partial Reset mode discussed in Section 4.2.1, these correspond to two distinct optimization objectives, namely “success_at_end” and “success_once”. Our codebase supports both objectives. We refer to the latter as the Valid Action Mask setting, which can be enabled by setting “auto_reset=False” and “ignore_terminations=False”. Figure 5(c) visualizes this option. In experiments, we observed that under the GRPO setting, the valid action mask objective is generally more effective for improving policy performance. A detailed analysis is provided in the following section.
Loss Normalization by Trajectory Length
Unlike in LLM tasks, where longer responses are often more likely to be correct, successful trajectories in robotics are typically not longer than failed ones. To ensure that both successful and failed trajectories contribute equally during optimization, we normalize the policy loss by the trajectory length under the Valid Action Mask setting. Specifically, if a trajectory $\tau_{i}$ has $T_{i}^{\text{succ}}$ valid timesteps, the contribution of each timestep to the objective is scaled by $1/T_{i}^{\text{succ}}$. The GRPO objective under this setting is:
@@ 4.2.2 GRPO
Success Rate Filter
Inspired by the dynamic sampling strategy in DAPO (Yu et al., 2025b), we introduce a success-rate filter for GRPO. This filter discards groups in which all trajectories either succeed or fail [^3] Following the practice in VeRL (Sheng et al., 2025), we set both lower and upper bounds on the average return of all trajectories within a group., since the advantage is non-zero only when both successful and failed trajectories coexist. Empirically, we find that this filtering mechanism accelerates convergence and consistently improves policy performance.
5 Experiment Results
In this section, we systematically investigate four key questions concerning the effectiveness of the proposed RLinf-VLA framework:
(1) Is RLinf-VLA high-performance? We evaluate RLinf-VLA on two representative testbeds, LIBERO and ManiSkill. The results demonstrate that RLinf-VLA achieves a success rate of 96.8% in large-scale mixed-task training involving up to 130 tasks, highlighting its strong capability to support large-scale multi-task learning.
(2) Is RLinf-VLA high-efficiency? We benchmark the framework across both GPU-parallelized and CPU-parallelized simulators, and observe that the optimal configuration differs across simulator types. This finding underscores the necessity of supporting diverse allocation modes. Notably, RLinf-VLA achieves up to $1.88\times$ speedup over existing frameworks.
(3) What are the best practices for applying PPO and GRPO to VLA training? Through extensive ablation studies, we identify the key factors that govern training performance. Moreover, we provide insights into the selection of common hyperparameters, offering practical guidelines for effectively deploying RL in VLA settings.
(4) Can models trained in simulation be deployed in the real world? To assess sim-to-real transfer, we conduct preliminary evaluations on six unseen object pick-and-place tasks. The results show that RL-trained models achieve higher zero-shot deployment success rates compared to SFT-trained models, demonstrating the promise of RL-based training for real-world applications.
@@ 5 Experiment Results
Evaluation. Unless otherwise specified, all curves are smoothed using a Gaussian filter ($\sigma=1$), and success rate is computed under the “success_once” criterion, where an episode is considered successful if the success state is reached at least once.
5.1 High-performance
5.1.1 ManiSkill
Experiments Setup
  1. 1.
    Training tasks: “PutOnPlateInScene25Main-v3”. As previously mentioned, it is adopted in RL4VLA (Liu et al., 2025).
  2. 2.
    Evaluation tasks: Various out-of-distribution (OOD) test sets on vision, semantic, and position. Also adopted from RL4VLA. This part checks the capacity of VLA models on vision, language, and action, respectively. All models are evaluated under each sub-setting using 256 randomly episodes.
  3. 3.
    Hyperparameters: Please refer to YAML files in RLinf-VLA open-sourced code, including maniskill_ppo_openvla.yaml, maniskill_ppo_openvlaoft.yaml, maniskill_grpo_openvla.yaml, and maniskill_grpo_openvlaoft.yaml.
  4. 4.
@@ 5.1.1 ManiSkill
    Base models: For the OpenVLA model, we adopt the pre-trained checkpoint available at HuggingFace [^4] https://huggingface.co/gen-robot/openvla-7b-rlvla-warmup. For the OpenVLA-OFT model, we perform our own LoRA fine-tuning using motion planning data collected from the “PutOnPlateInScene25Main-v3” task. The resulting LoRA model weights are also provided at HuggingFace [^5] https://huggingface.co/RLinf/RLinf-OpenVLAOFT-ManiSkill-Base-Lora.
Training Results
Figure 6 reports the training performance on ManiSkill across 25 tasks, using the OpenVLA and OpenVLA-OFT models trained with PPO and GRPO. Figure 6(a) shows the results for OpenVLA, while Figure 6(b) presents the OpenVLA-OFT results. Across all settings, reinforcement learning provides substantial performance gains, improving success rates by 45%–70% compared to the baseline. Notably, PPO consistently outperforms GRPO and exhibits greater stability for both OpenVLA and OpenVLA-OFT in the 25-task ManiSkill setting.
(a) OpenVLA
(b) OpenVLA-OFT
Figure 6: Training curves on ManiSkill “PutOnPlateInScene25Mani-v3” with OpenVLA and OpenVLA-OFT models, using PPO and GRPO algorithms. PPO consistently outperforms GRPO and exhibits greater stability.
Evaluation Results
Table 2 summarizes the evaluation results of RLinf-VLA models. For OpenVLA, we include four variants: the base model (“OpenVLA (Base)”), the RL4VLA baseline trained with PPO (“RL4VLA (PPO)”), and our RLinf-VLA models trained with GRPO (“OpenVLA (RLinf-GRPO)”) and PPO (“OpenVLA (RLinf-PPO)”). For OpenVLA-OFT, we report results on three variants: the LoRA fine-tuned base model (“OpenVLA-OFT (Base)”), along with the RLinf-VLA models trained using GRPO (“OpenVLA-OFT (RLinf-GRPO)”) and PPO (“OpenVLA-OFT (RLinf-PPO)”).
Although there is no fundamental algorithmic difference between the PPO-trained OpenVLA models in RLinf-VLA and in RL4VLA, our framework is carefully optimized to make better use of GPU resources. This enables training for more steps within a feasible time budget, leading to slightly better performance in RLinf-VLA compared to RL4VLA.
It is important to note that a direct comparison between OpenVLA and OpenVLA-OFT on OOD evaluation is not entirely fair, since their base models differ substantially in performance. Specifically, while OpenVLA improves from $39.10\%$ to $81.93\%$ after RL fine-tuning and OpenVLA-OFT improves from $18.29\%$ to $77.05\%$, the relative gain of OpenVLA-OFT is slightly higher despite its much weaker starting point. This indicates that the base model is very important and plays a crucial role in the final generalization.
Table 2: Evaluation results on ManiSkill. Values denote success rates.
In-Distribution Out-Of-Distribution
Vision Semantic Execution Avg.
OpenVLA (Base) 53.91% 38.75% 35.94% 42.11% 39.10%
RL4VLA (PPO) 93.75% 80.47% 75.00% 81.77% 79.15%
OpenVLA (RLinf-GRPO) 84.38% 74.69% 72.99% 77.86% 75.15%
OpenVLA (RLinf-PPO) 96.09% 82.03% 78.35% 85.42% 81.93%
OpenVLA-OFT (Base) 28.13% 27.73% 12.95% 11.72% 18.29%
OpenVLA-OFT (RLinf-GRPO) 94.14% 84.69% 45.54% 44.66% 60.64%
OpenVLA-OFT (RLinf-PPO) 97.66% 92.11% 64.84% 73.57% 77.05%
5.1.2 LIBERO
Experiments Setup
  1. 1.
    Tasks: We train a single unified model on all 130 tasks in LIBERO and evaluate its performance across the five LIBERO task suites: LIBERO-Object, LIBERO-Spatial, LIBERO-Goal, LIBERO-10, and LIBERO-90.
  2. 2.
    Hyperparameters: See the YAML configuration file in the RLinf-VLA open-source repository: libero_130_grpo_openvlaoft.yaml.
  3. 3.
@@ 5.1.2 LIBERO
    Base models: We fine-tune OpenVLA-OFT with LoRA ($r=32$) using demonstrations from all five LIBERO task suites: LIBERO-Object, LIBERO-Spatial, LIBERO-Goal, LIBERO-10, and LIBERO-90. The resulting LoRA model weights are also provided at HuggingFace  [^6] https://huggingface.co/RLinf/RLinf-OpenVLAOFT-LIBERO-130-Base-Lora.
Training Results
Figure 7 illustrates the training curve of OpenVLA-OFT with the GRPO algorithm across the five task groups. The x-axis represents the number of training epochs, while the y-axis indicates the corresponding success rate. The results show that the success rate improves substantially from approximately $73\%$ to $98\%$, yielding an overall performance gain of about $30\%$. These findings highlight the effectiveness of the GRPO design in RLinf-VLA for enhancing multi-task training.
Figure 7: Training curve of OpenVLA-OFT with GRPO on LIBERO-130.
Evaluation Results
Table 3 summarizes the evaluation results of the OpenVLA-OFT model trained on LIBERO-130 using RLinf-VLA across the five LIBERO task groups. For each task group, the success rate is computed as the average over 50 evaluation episodes per task. For instance, in the “LIBERO-90”, the reported success rate is averaged over a total of $90\times 50=4500$ episodes. The results show that the LIBERO-Spatial and LIBERO-Object group reaches over $99\%$, the LIBERO-Goal and LIBERO-90 task groups achieve up to $98\%$, and the LIBERO-10 task group exceeds $93\%$, yielding an overall average success rate of $98.11\%$. This corresponds to an average performance improvement of $32.68\%$ across task categories. Notably, our approach adopts a single model for all tasks, in contrast to the existing RL+VLA paradigm that trains and evaluates a separate model for each task group. This demonstrates RLinf-VLA’s capability to support large-scale multi-task RL.
Table 3: Evaluation results of the unified model on the five LIBERO task groups.
Spatial Object Goal 10 90 Avg.
OpenVLA-OFT (Base) 72.18% 71.48% 64.06% 48.44% 70.97% 65.43%
OpenVLA-OFT (RLinf-GRPO) 99.40% 99.80% 98.79% 93.95% 98.59% 98.11%
$\Delta$ Improvement +27.22{+27.22} +28.32{+28.32} +34.73{+34.73} +45.51{+45.51} +27.62{+27.62} +32.68{+32.68}
5.2 High-efficiency
To evaluate the efficiency of our framework, we conduct end-to-end experiments comparing throughput under different GPU allocation modes.
5.2.1 Experiment Setup
Evaluation Metric.
We use throughput as the evaluation metric for efficiency. Specifically, throughput is defined as the total number of rollout environment frames divided by the total wall-clock time of one training epoch, which approximately equals the sum of rollout time and training time. The evaluation is conducted on 8, 16, and 32 NVIDIA H100 (80GB) GPUs, distributed across 1, 2, and 4 nodes, respectively.
Tasks.
We evaluate our framework on both ManiSkill and LIBERO setups.
For ManiSkill, we select the “PutCarrotOnPlateInScene-v2” task for a quick test. It is adapted from “PutCarrotOnPlateInScene-v1” in ManiSkill (Tao et al., 2025), with modifications to the reset function. [^7] In the original reset implementation, additional simulation steps were applied to ensure that all objects remained static. However, this significantly increased the reset time compared to the step function and was incompatible with partial reset. In practice, we observed that these extra steps had a negligible impact on object states in the “PutCarrotOnPlateInScene” task, since the initial position of the carrot was already carefully calibrated. We therefore delete the additional simulation steps in the reset function. We use 256 parallel environments, each running for 80 steps.
For LIBERO, we adopt the LIBERO-10 task set for a quick test. The number of parallel environments is set to 64, 128, and 256 for 8-, 16-, and 32-GPU setups, respectively, with the corresponding number of environment steps set to 4096, 2048, and 1024. Since vectorized environments in LIBERO rely on multi-processing, the number of physical CPU cores per node becomes the upper bound for efficient rollout. Therefore, we scale the number of parallel environments in proportion to the number of nodes, and consequently to the total number of GPUs.
Baselines.
For LIBERO, we compare against SimpleVLA-RL (Li et al., 2025), an open-source framework for VLA+RL training built on VeRL. Since SimpleVLA-RL only supports the colocated mode, we use it as the baseline and additionally evaluate our colocated mode and hybrid mode (with one-stage and two-stage fine-grained pipelining). We omit the disaggregated mode results for LIBERO, as during training the Training component never reuses the same GPUs as the rollout process. Consequently, even when accounting for the offload–onload overhead, the hybrid (one-stage) configuration provides superior performance to the disaggregated mode.
GPU Allocation Modes.
Tables 4 and 5 summarize the GPU allocation and offloading configurations used in our experiments.
@@ 5.2.1 Experiment Setup
Table 4: The setting of enable_offload among different GPU allocation modes.
@@
Table 5: Hyperparameters of GPU allocation modes.
@@ 32 GPUs Colocated 0-31 0-31 0-31
(a) GPU-parallelized Simulator ManiSkill with OpenVLA
(b) GPU-parallelized Simulator ManiSkill with OpenVLA-OFT
(c) CPU-parallelized Simulator LIBERO with OpenVLA-OFT
Figure 8: Throughput for different settings. “pipe” denotes the “pipeline_stage_num” for fine-grained pipelining.
5.2.2 Results
Figure 8 presents the end-to-end throughput of RLinf and baseline methods in VLA+RL training under different cluster sizes and placement strategies.
As shown in Figure 8(a), for the GPU-parallelized ManiSkill simulator with OpenVLA, RLinf-VLA achieves a $1.88\times$ speedup compared to the baseline disaggregated mode when using the hybrid configuration with pipe=2 on 8 GPUs. When scaling to 16 and 32 GPUs, additional overhead arises from model loading, offloading, and state switching. Consequently, the colocated mode generalizes less effectively than the disaggregated mode. However, the hybrid (pipe=2) configuration mitigates these costs through pipelined execution that overlaps computation and communication, still achieving a $1.61\times$$1.69\times$ speedup over the disaggregated mode. Furthermore, comparing hybrid configurations with pipe=1 and pipe=2, we observe that deeper pipelines yield better performance, suggesting that fine-grained pipelining effectively reduces GPU idle time (“bubbles”) during the rollout phase.
As shown in Figure 8(b), the results for OpenVLA-OFT exhibit substantial differences. In this case, the colocated mode and the hybrid (pipe=1) mode outperform the disaggregated mode and the hybrid (pipe=2) mode. This discrepancy stems from the use of action chunks. Since OpenVLA-OFT predicts an entire chunk of actions in each inference step while the simulator executes them sequentially, the simulator’s step execution time becomes much longer than the generation time (shifting from a 1:1 ratio to approximately 15:1). Consequently, the benefit of pipelining to hide generation latency diminishes. Moreover, since ManiSkill is GPU-parallelized and the rollout speed scales linearly with the number of parallel environments, allocating more GPU resources to the simulator makes the colocated and hybrid (pipe=1) modes more favorable.
Notably, when comparing the colocated mode and the hybrid (pipe=1) mode, we observe that the latter achieves progressively better throughput as the number of GPUs increases. This is because, in the colocated configuration, each GPU hosts both the simulator and the actor model simultaneously  [^8] Note that in the standard colocated setting, both the simulator and the actor model fully occupy all GPUs, and context switching is handled via onload and offload. However, since ManiSkill is GPU-parallelized, this approach incurs excessive overhead during multi-step interactions. Therefore, in our implementation, each GPU allocates half of its memory to each component.. In contrast, the hybrid configuration splits the cluster evenly, with each half of the GPUs dedicated to one component. As a result, the communication overhead in the colocated mode becomes increasingly non-negligible as the number of GPUs grows.
Figure 9: Latency breakdown for LIBERO with the OpenVLA-OFT setting.
For the CPU-parallelized LIBERO simulator with OpenVLA-OFT [^9] Previous results have shown that the disaggregated mode consistently underperforms the hybrid mode, as the training process in the disaggregated setup cannot utilize all GPUs. Therefore, the disaggregated mode is excluded from comparison here., the trend differs once again. As shown in Figure 8(c), the colocated mode achieves the best performance, yielding a $1.34\times$$2.27\times$ speedup over the SimpleVLA-RL implementation.
We further analyze the source of this speedup in Figure 9, where “Rollout” represents the total time for multi-step interactions between the generator and the simulator. Performance improvements are observed in both the rollout and training stages. Specifically, during rollout, RLinf-VLA employs a vectorized environment implementation that is more efficient than the hand-written multi-process environment workers used in SimpleVLA-RL. In addition, our implementation avoids redundant log-probability recomputation present in SimpleVLA-RL, further improving the rollout performance. During the training stage, RLinf adopts several system-level optimizations, including an adaptive communication mechanism, which leads to greater efficiency gains as the number of GPUs increases. Compared with the colocated mode, the hybrid mode incurs longer rollout times since LIBERO is a CPU-parallelized simulator. In this case, computation overlap provides little benefit, and allocating more compute resources to generation yields higher returns.
In summary, different types of simulators require distinct allocation modes, and varying interaction characteristics further necessitate adaptive configurations. RLinf-VLA supports three flexible allocation modes, enabling optimal system configurations under diverse scenarios.
5.3 Ablation Study
In this section, we present ablation studies under different setups and summarize key insights for PPO and GRPO training. In Section 5.3.1, we highlight PPO-specific strategies, such as step-level value estimation and partial reset. In Section 5.3.2, we discuss GRPO techniques, including trajectory length normalization, valid action masking, and success filtering. Finally, in Section 5.3.3, we analyze the effects of rollout data size and LoRA-based adaptations.
5.3.1 Tips for PPO
(a) Success rate in ManiSkill.
(b) Value loss in ManiSkill.
(c) Success rate in LIBERO-Goal
Figure 10: Ablation of value estimation type on PPO with action chunks.
(a) OpenVLA
(b) OpenVLA-OFT
Figure 11: Ablation study on partial reset with PPO in ManiSkill.
Action-level value estimation outperforms chunk-level estimation for PPO with action chunks.
Figure 10 shows the PPO training curves comparing action-level and chunk-level value estimation as described in Section 4.2.1. The x-axis denotes the number of training epochs, while the y-axis reports the success rate (Figure 10(a)) and value loss (Figure 10(b)) on the ManiSkill “PutOnPlateInScene25Main” task using the OpenVLA-OFT model. Step-level estimation consistently yields higher success rates and lower value loss, demonstrating more effective learning and faster policy improvement. This advantage is not task-specific: similar results are observed in the LIBERO-Goal benchmark (Figure 10(c)).
Partial reset substantially improves sample efficiency.
Figure 11 presents the PPO training curves for the Partial Reset and Fixed Episode Length rollout modes discussed in Section 4.2.1. The x-axis indicates the number of training epochs, and the y-axis shows the success rate. Since the optimization objective is “success_once,” Partial Reset leads to a significantly higher success rate. For a given number of training epochs, the success rate under Partial Reset consistently exceeds that of the Fixed Episode Length mode. This trend is observed regardless of the model type (OpenVLA or OpenVLA-OFT).
5.3.2 Tips for GRPO
Trajectory length normalization in GRPO.
Figure 12: Ablation study on trajectory length normalization.
As discussed in Section 4.2.2, normalizing the loss by trajectory length aims to reduce bias when episodes vary in length. Figure 12 reports results on LIBERO-Goal, where the x-axis denotes the number of training epochs and the y-axis the success rate. The results show that incorporating trajectory length normalization (“w/ Norm”) can lead to substantially higher performance compared with the unnormalized setting (“w/o Norm”).
Valid action mask in GRPO.
We also study the effect of valid action masking, introduced in Section 4.2.2. Since the optimization objective is defined with respect to “success_once”, excluding actions after reaching the success state improves sample efficiency and avoids redundant updates. Moreover, applying the valid action mask naturally results in shorter trajectories, which further benefits trajectory length normalization. Figure 13(a) shows results on LIBERO-Goal. Comparing the “w/o Mask, w/o Norm” and “w/ Mask, w/o Norm” curves, the setting with a valid action mask achieves consistently better performance. The “w/ Mask, w/ Norm” curve demonstrates that combining the mask with trajectory length normalization provides an additional improvement.
(a) LIBERO-Goal
(b) ManiSkill
Figure 13: Ablation studies on valid action mask in GRPO.
However, the effect of these techniques is task-dependent. In the ManiSkill setting, we do not observe clear benefits from either valid action masking or trajectory length normalization. As shown in Figure 13(b), the performance of the “w/ Mask, w/ Norm” configuration is comparable to that of the “w/o Mask, w/o Norm” baseline, suggesting that the effectiveness of these strategies may vary across environments.

### v2

@@ VIII-C1 Details for Multi-Granularity Support
VIII-C2 Details for Loss Normalization by Trajectory Length
As mentioned in the main text, we normalize the policy loss by the trajectory length under the Valid Action Mask setting. Specifically, if a trajectory $\tau_{i}$ has $T_{i}^{\text{succ}}$ valid timesteps, the contribution of each timestep to the objective is scaled by $1/T_{i}^{\text{succ}}$. The GRPO objective under this setting is:
@@ VIII-C2 Details for Loss Normalization by Trajectory Length
IX Experiment Setup
TABLE III: Hyperparameter settings for training.
Parameter OpenVLA [18] OpenVLA-OFT [17] OpenVLA-OFT OpenVLA-OFT
ManiSkill [39] ManiSkill LIBERO [21] RoboTwin [8]
PPO GRPO PPO GRPO GRPO GRPO
# Parallel Envs 128 256 128 256 64 128
Max Episode Steps 80 80 80 80 512 200a
Max Steps per Rollout Epoch 160 80 160 80 512 200a
Group Size 1 8 1 8 8 8
Rollout Epoch 1 1 1 1 64 8
Partial Reset True False True False False False
Valid Action Mask False True False True True True
Action Chunk Size 1 1 8 8 8 25
Global Batch Size 640 640 640 640 16384 1024
Micro Batch Size 40 40 40 40 32 32
is_lora [13] True True True True True
Learning Rate 1e-4 1e-5 1e-4 1e-4 2e-5 1e-4
adam_eps 1e-8 1e-8 1e-8 1e-8 1e-8 1e-5
Clip Ratio ($\epsilon$) (0.2, 0.28) (0.2, 0.28) (0.2, 0.28) (0.2, 0.28) (0.2, 0.28) (0.2, 0.28)
$\gamma$ 0.99 0.99
GAE $\lambda$ 0.95 0.95
Temperature (train) 1.0 1.0 1.0 1.0 1.6 1.6
  • a
    Note that the “pick dual bottles” and “handover block” task in RoboTwin requires 100 and 400 steps respectively.
IX-A Details for Performance Experiments
@@
This part is the details of training and evaluation for High-Performance experiments, including metric for evaluation, base models selection and hyperparameter settings for training. Unless otherwise specified, all curves are smoothed using a Gaussian filter ($\sigma=1$), and success rate is computed under the “success_once” criterion, where an episode is considered successful if the success state is reached at least once.
IX-A1 Base Models for Training
@@
For the OpenVLA (Base) in ManiSkill, we adopt the pre-trained checkpoint OpenVLA (Base) from RL4VLA [22]. For the OpenVLA-OFT (Base) in ManiSkill, we perform our own LoRA fine-tuning to get OpenVLA-OFT (Base) using motion planning data collected from the “PutOnPlateInScene25Main-v3” task.
@@
For OpenVLA-OFT (Base) in LIBERO 130 tasks, we fine-tune OpenVLA-OFT with LoRA using demonstrations from all five task suites in LIBERO.
For OpenVLA-OFT (Base) in RoboTwin, we adopt the pre-trained checkpoint OpenVLA-OFT (Base) from SimpleVLA-RL [20]. Besides, we also reuse the train/eval seeds from SimpleVLA-RL [20].
IX-A2 Hyperparameter Settings for Training
In Table III, we list the specific hyperparameters used for different experiments.
IX-A3 Metric for Evaluation
For supervised fine-tuned models, we set do_sample=False during evaluation. And for RL-trained models, we set do_sample=True during evaluation, and conduct three times evaluation for each model. And we report the mean and standard deviation of the success rates. (Now only the LIBERO setting has standard deviation.)
IX-B Details for Efficiency Experiments
@@
TABLE IV: The setting of enable_offload among different GPU allocation modes.
@@
TABLE V: Hyperparameters of GPU allocation modes.
@@ 32 GPUs Colocated 0-31 0-31 0-31
IX-B1 Hyperparameters for Efficiency Experiments
See Table IV for how we enable the offload configurations and Table V for how we configure the GPU allocation.
IX-B2 Metric for Efficiency
We use throughput as the evaluation metric for efficiency. Specifically, throughput is defined as the total number of rollout environment frames divided by the total wall-clock time of one training epoch, which approximately equals the sum of rollout time and training time.
IX-B3 Tasks for Efficiency
We evaluate our framework on all tasks in ManiSkill, LIBERO and RoboTwin.
For ManiSkill, we select the “PutCarrotOnPlateInScene-v2” task for a quick test. It is adapted from “PutCarrotOnPlateInScene-v1” in ManiSkill, with modifications to the reset function. In the original reset implementation, additional simulation steps were applied to ensure that all objects remained static. However, this significantly increased the reset time compared to the step function and was incompatible with partial reset. In practice, we observed that these extra steps had a negligible impact on object states in the “PutCarrotOnPlateInScene” task, since the initial position of the carrot was already carefully calibrated. We therefore delete the additional simulation steps in the reset function. The evaluation is conducted on 8, 16, and 32 NVIDIA H100 (80GB) GPUs, and we use 256 parallel environments, each running for 80 steps.
For LIBERO, we adopt the LIBERO-Long task set for a quick test, which is the longest task set in LIBERO. The number of parallel environments is set to 64, 128, and 256 for 8-, 16-, and 32-GPU setups, respectively, with the corresponding number of environment steps set to 4096, 2048, and 1024.333Since vectorized environments in LIBERO rely on multi-processing, the number of physical CPU cores per node becomes the upper bound for efficient rollout. Therefore, we scale the number of parallel environments in proportion to the number of nodes, and consequently to the total number of GPUs.
For RoboTwin, we select the place_empty_cup task for a quick test, which requires the robot to use an arm to place the empty cup on the coaster. The number of parallel environments is set to 64, 128, and 256 for 8-, 16-, and 32-GPU setups, respectively, with the corresponding number of environment steps set to 800, 400, and 200. We have the same GPU allocation setting as LIBERO-Long, due to the GPU resources being the upper bound for the rollout.
IX-B4 Baselines for Efficiency Experiments
For LIBERO and RoboTwin, we compare against SimpleVLA-RL [20], an open-source framework for RL of VLA Models training built on VeRL. Since SimpleVLA-RL only supports the colocated mode, we use it as the baseline and additionally evaluate our colocated mode and hybrid mode (with one-stage and two-stage fine-grained pipelining). We omit the disaggregated mode results for LIBERO and RoboTwin, as during training the Training component never reuses the same GPUs as the rollout process. Consequently, even when accounting for the offload–onload overhead, the hybrid (one-stage) configuration provides superior performance to the disaggregated mode.
X Additional Experimental Results
X-A Evaluation of OOD for ManiSkill
Table VI is the detailed results ManiSkill’s experiments.
TABLE VI: Detailed breakdown of Out-Of-Distribution evaluation results on ManiSkill. Values denote success rates (%).
Out-Of-Distribution
Method Vision Semantic Execution
OpenVLA (Base) 38.75 35.94 42.11
OpenVLA (RLinf-GRPO) 74.69 72.99 77.86
OpenVLA (RLinf-PPO) 82.03 78.35 85.42
OpenVLA-OFT (Base) 27.73 12.95 11.72
OpenVLA-OFT (RLinf-GRPO) 84.69 45.54 44.66
OpenVLA-OFT (RLinf-PPO) 92.11 64.84 73.57
X-B Ablation Studies